D Question 5 1 pts A tailor must make 4 of the same size suits for a client. The first suit took 12.5 hours to make and the last suit took 8.6 hours to make, what was the tailors learning curve % of this run of suits assuming a steady state of production is never reached? Express your answer in % to the nearest 0.1%
Answer #1

Time taken for first suit=12.5

Time taken for 4th suit=8.6

Fraction of Time reduced=8.6/12.5=0.688 i.e 68.8% of time is taken for 4th suit.

Thus (1+r)^t=0.688

Thus r=-0.09 where r is learning coefficient

Thus 9% less time is take everyear which makes learning curve %=9%